OpenStudy (anonymous):

[2.02] Solve: 4x − 10 + 16 = 22 4 12 7 −4

OpenStudy (yacoub1993):

\[4x-10+16=22\] \[4x=22+10-16\] \[4x=16\] \[x=\frac{ 16 }{ 4 }\] x=4

OpenStudy (yacoub1993):

so u go with the frist choice "A"

OpenStudy (yacoub1993):

Good luck